Question 3

Presented below are selected transactions at Ayayai Company for 2017.

Jan. 1 Retired a piece of machinery that was purchased on January 1, 2007. The machine cost £59,570 on that date. It had a useful life of 10 years with no residual value.
June 30 Sold a computer that was purchased on January 1, 2014. The computer cost £35,540. It had a useful life of 5 years with no residual value. The computer was sold for £13,650.
Dec. 31 Discarded a delivery truck that was purchased on January 1, 2013. The truck cost £37,850. It was depreciated based on a 6-year useful life with a £3,710 residual value.


Journalize all entries required on the above dates, including entries to update depreciation, where applicable, on assets disposed of. Ayayai Company uses straight-line depreciation. (Assume depreciation is up to date as of December 31, 2016.) (Credit account titles are automatically indented when amount is entered. Do not indent manually.)

Date

Account Titles and Explanation

Debit

Credit

Jan. 1June 30Dec. 31

Jan. 1June 30Dec. 31

(To record depreciation)

June. 30

(To record the sale of equipment)

Jan. 1June 30Dec. 31

(To record depreciation)

Dec. 31

(To record the discarding of equipment)
Answer #1

(1) machinery purchased in 2007 with 10 years useful life will be fully depreciated by 2016 December.

there will be £59,570 in accumulated depreciation so we will reverse it.

Date

Account Titles and Explanation

Debit

Credit

Jan. 1

Accumulated depreciation-Equipment

£59,570

Equipment

£59,570

(2) we will find book value on June 30 after giving effect of depreciation. half year depreciation from January to June

Date

Account Titles and Explanation

Debit

Credit

Jan. 30

Depreciation expense

£3,554

Accumulated Depreciation - computer

£3,554

£35,540/5years * 1/2 year

Debit Credit

June. 30

cash

£13,650

Accumulated depreciation-Equipment

24,878

Gain on disposal of equipment

2,988

Equipment

£35,540

(To record the sale of equipment)

depreciation accumulated from January 2014 upto June30 2017

£35,540 / 5years = 7,108 for 2014 + 7108 for 2015 + 7108 for 2016 + 3554 for 2017 = 24,878

Gain = purchase cost - depreciation - selling price

=35540-24878-13650

=2,988

(3) straight line depreciation = cost-residual value/useful life

=37850-3710/6

=5690

Debit Credit

Dec. 31

Depreciation expense

£5,690

Acummulated depreciation

£5,690

(To record depreciation)

accumulated depreciation from January 2013 to Dec 31 2017

=5690 for 2013+5690 for 2014 +5690 for 2015+5690 for 2016+5690 for 2017

=28450

Debit Credit

Dec. 31

Loss on disposal of Truck/Asset

£9,400

Accumulated depreciation

£28,450

Equipment

£37,850

(To record the discarding of equipment)